# BPMN Assistant: An LLM-Based Approach to Business Process Modeling

### v1

@@ 4.1 Graph Edit Distance (GED) and Relative Graph Edit Distance (RGED)
GED quantifies the cost of transforming one BPMN diagram into another by counting the minimum number of graph edit operations (e.g., node insertion, deletion, or substitution) required. It is a fundamental metric for evaluating the accuracy of diagram generation and modifications.
RGED normalizes the GED by considering the complexity of the involved graphs. This metric allows for a relative comparison of diagram similarity, regardless of their absolute sizes, providing a more interpretable evaluation score.
$\text{Relative GED}=\frac{\text{GED}(G_{1},G_{2})}{\text{GED}(G_{1},\text{Empty})+\text{GED}(G_{2},\text{Empty})}$ (1)
@@
To implement GED, we utilized the NetworkX Python library, which provides efficient algorithms for computing graph edit distance. Unlike prior approaches that rely on domain-specific heuristics or ML-based enhancements [35], our implementation applies standard GED operations to measure similarity between BPMN diagrams.
While our evaluation focuses on structural similarity using GED and RGED, these metrics inherently reflect aspects of usability and clarity when compared against human-generated BPMN models. However, they do not explicitly capture higher-level qualitative factors such as readability, model simplicity, or semantic coherence, which are emphasized in studies like those by Huang and Kumar [38] and Pavlicek et al. [39].
@@ 4.2 Comparison with Baseline XML Editing
To establish the effectiveness of BPMN Assistant, we conducted a comparative analysis between our JSON-based BPMN representation and direct XML manipulation through language model prompting. The baseline method involves prompting the model to directly generate or modify BPMN XML specifications, which are then rendered into diagrams.
For this comparison, we selected a representative set of diagram modification tasks and executed them using both representations. We evaluated the approaches using RGED and its derived similarity score, as defined in Equations (1) and (2).
@@ 4.3 Generation Accuracy
The comparative analysis revealed that our JSON-based representation achieved an average similarity score of 0.70 compared to 0.69 for direct XML generation. While this indicates a slight numerical advantage for JSON, the difference is minimal, suggesting that both representations perform similarly in terms of structural accuracy. However, JSON demonstrated greater reliability, with fewer total failures across models.
@@
Beyond similarity scores, JSON also outperformed XML in terms of efficiency. As shown in Table 7 and illustrated in Figure 3, JSON-based BPMN generation was significantly faster, with a mean latency of 14.41 seconds compared to 24.07 seconds for XML. Additionally, as depicted in Figure 4, JSON required more input tokens (2604.34 vs. 466.79 on average) but produced more concise outputs (849.29 vs. 1834.55 tokens). This trade-off suggests that JSON requires a richer input prompt but benefits from reduced processing time and smaller outputs, making it more efficient overall.
@@
Mean Latency (seconds) 14.41 24.07
Average Input Tokens 2604.34 466.79
Average Output Tokens 849.29 1834.55
@@ 4.4 Editing Capabilities
In addition to diagram generation, we evaluated the models’ ability to interpret natural language editing instructions and perform the requested modifications. For this evaluation, we used a binary success/fail metric, as editing tasks involve correctly understanding the request and applying appropriate modifications to an existing diagram.
To determine validity, we implemented a Python-based validation function that checked whether the generated BPMN structure adhered to syntactic constraints. If a modification resulted in an invalid BPMN diagram, it was automatically marked as a failure. For valid outputs, an additional verification step assessed whether the requested edit had been correctly applied. This evaluation involved both automated checks and manual review to ensure accuracy.
The results, summarized in Table 8, show that models consistently achieved higher success rates when using our JSON-based approach compared to direct XML manipulation. JSON-based edits were more reliably executed across all tested models, likely due to the structured nature of JSON, which facilitates clearer representation of modifications.
@@
Figure 5 provides a visual representation of these results, reinforcing the observed performance gap between JSON and XML approaches.
@@
Our performance analysis also revealed substantial differences between JSON and XML approaches across both latency and token utilization metrics. As shown in Table 9, the JSON-based approach demonstrated significantly faster average processing times, requiring 21.46 seconds compared to 46.98 seconds for XML generation.
Token usage patterns showed an inverse relationship between input and output requirements for the two approaches. JSON-based generation demanded higher input token counts, averaging 19,567.70 tokens compared to XML’s 5,102.74 tokens. However, this was offset by more compact outputs, with JSON generating an average of 364.14 tokens compared to XML’s 2,665.32 tokens.
@@
Average Latency (s) 21.46 46.98
Average Input Tokens 19,567.70 5,102.74
Average Output Tokens 364.14 2,665.32
@@
These performance characteristics suggest a clear trade-off between the approaches: while JSON requires more substantial input context, it achieves faster processing times and produces more concise outputs. The higher input token requirement of the JSON approach is offset by its reduced latency and more manageable output sizes, potentially offering better responsiveness in real-world BPMN editing scenarios. When combined with the higher editing success rates previously discussed, these performance metrics further support the advantages of the JSON-based approach for practical BPMN manipulation tasks.
@@ 5 Discussion
Our evaluation demonstrates that the JSON-based representation consistently outperforms direct XML manipulation across multiple metrics. Higher success rates in editing tasks (Table 8) suggest that structured intermediate representations are particularly valuable for complex model modifications. Performance differences between models indicate that model selection significantly impacts system effectiveness, with Claude 3.5 Sonnet and o3-mini showing particularly strong performance.
The inverse relationship between input and output token usage in the JSON versus XML approaches (Table 9) reveals interesting trade-offs that likely favor the JSON approach. Although the JSON approach requires more context in the prompts, it produces more concise output and faster processing times. This combination is particularly advantageous since input tokens typically cost less than output tokens, making the JSON approach potentially more cost-effective despite using more prompt tokens.
Importantly, BPMN Assistant is designed to make business process modeling more accessible to users without a technical background. Instead of requiring knowledge of specialized diagramming tools or technical formats like XML, users can simply describe a process in natural language. The system then automatically generates a professional diagram based on this description. This reduces the need for technical expertise, allowing business professionals to contribute directly to process design and updates, which can lead to more accurate and timely documentation.
The system’s ability to handle natural language editing commands has significant implications for process modeling workflows. The high success rates for editing tasks suggest that BPMN Assistant could reduce the technical barriers to process model maintenance, potentially leading to more accurate and up-to-date process documentation in practice.
Unlike previous tools such as BPMN-Chatbot [19] and ProMoAI [20], our system emphasizes model modification capabilities in addition to generation. A structured JSON representation offers advantages over both direct XML manipulation and alternative graph representations, such as the Graphviz DOT language  [^1] https://graphviz.org/doc/info/lang.html, particularly in maintaining BPMN standard compliance and supporting complex editing operations.
Although our approach shows promise, several challenges remain. The dependency on LLM quality affects system reliability, and the current implementation’s limited support for BPMN elements restricts its application in complex scenarios. Additionally, the higher input token requirements of our JSON-based approach may impact costs in high-volume applications.
@@ 6 Limitations and Future Work
BPMN Assistant, while demonstrating promising results, has several limitations that should be acknowledged. A key limitation is the support for only a subset of BPMN elements (general tasks, user tasks, service tasks, exclusive gateways, parallel gateways, start events, and end events), which restricts the complexity of processes that can be modeled. However, since the focus of this work is on executable process models rather than collaboration diagrams, the absence of pools and lanes does not significantly impact our primary research objectives.
@@
The performance of BPMN Assistant is heavily dependent on the underlying LLM. Our evaluation revealed significant variations in performance across different models, with Claude 3.5 Sonnet and o3-mini showing particularly strong results. Further investigation into the specific capabilities and limitations of different LLMs in the context of process modeling would provide valuable insights for future improvements.
@@ 7 Conclusion
This paper presented BPMN Assistant, a novel approach to business process modeling that leverages Large Language Models to bridge the gap between natural language descriptions and formal BPMN representations. Through comprehensive evaluation, we demonstrated that our JSON-based representation for BPMN diagrams outperforms direct XML manipulation across multiple key metrics. In generation tasks, the JSON approach achieved lower latency with an average processing time of 14.41 seconds compared to 24.07 seconds for XML generation, while maintaining comparable model quality with an average similarity score of 0.70.
The evaluation results particularly highlight the effectiveness of our JSON-based representation in editing tasks, where it achieved consistently higher success rates across all tested models compared to XML-based editing, with editing operations taking an average of 21.46 seconds compared to 46.98 seconds for XML manipulation. While the JSON approach requires higher input token counts, this is offset by significantly more compact outputs, averaging 364.14 tokens compared to XML’s 2,665.32 tokens. These findings suggest that structured intermediate representations can enhance the reliability and efficiency of LLM-based process modeling tools.
Despite these achievements, our work also reveals important areas for future research. The current limitations in supporting only a subset of BPMN elements indicate opportunities for expanding the system’s capabilities, particularly for modeling more complex executable processes. Additionally, the varying performance across different LLMs suggests potential benefits from model-specific optimization approaches.
In broader terms, this research represents a significant step toward making business process modeling more accessible to a wider audience. By enabling natural language interaction with process modeling tools, BPMN Assistant reduces the technical knowledge required to create and modify business process diagrams. This accessibility has potential implications beyond technical efficiency gains, as it could foster better communication between technical and non-technical stakeholders in organizations. When business experts can directly interact with modeling tools using everyday language, they may become more engaged in the process modeling activities, potentially leading to more accurate process documentation and improved alignment between business needs and implemented processes. The natural language approach also opens possibilities for process modeling in educational contexts, where students can focus on understanding process concepts rather than mastering technical modeling interfaces.
Looking ahead, future work could focus on expanding support for additional BPMN elements to enable modeling of more complex executable processes. The significant performance variations observed across different LLMs also suggest opportunities for developing specialized prompt engineering techniques for process modeling tasks. These improvements would further enhance the practical applicability of LLM-based process modeling tools in enterprise settings.
@@ Appendix A Appendix: Process Example
The following is an example process modeled using the BPMN Assistant based on the textual description:
@@
"The manager sends the mail to the supplier and prepares the documents. At the same time, the customer searches for the goods and picks up the goods."
@@
            "type": "parallelGateway",
            "id": "parallel1",
            "branches": [
@@
                        "label": "Send mail to supplier",
                        "type": "task",
                        "id": "task2",
                        "label": "Prepare the documents",
                        "type": "task",
                        "id": "task3",
                        "label": "Search for the goods",
                        "type": "task",
                        "id": "task4",
                        "label": "Pick up the goods",

### v2

@@ 4.2 Graph Edit Distance (GED) and Relative Graph Edit Distance (RGED)
GED quantifies the cost of transforming one BPMN diagram into another by counting the minimum number of graph edit operations (node insertion, deletion, or substitution) required. To address the semantic variability inherent in natural language generation (e.g., an LLM generating "Process Order" versus "Handle Order"), we implemented a custom two-stage evaluation pipeline [^2] Implementation available at: https://github.com/jtlicardo/bpmn-ged.
First, we perform semantic label normalization. Raw BPMN XML files are parsed into a graph structure, and a lightweight LLM (GPT-5 mini) is employed to map semantically equivalent labels to identical abstract tokens (e.g., mapping both "Submit Order" and "Send Order" to token "A"). This ensures that the metric evaluates logical flow accuracy rather than penalizing surface-level lexical differences.
Second, we compute the GED using the NetworkX library with a specific cost configuration designed to distinguish between minor syntactic errors and major structural hallucinations. The cost functions are defined as follows:
Insertion and Deletion
The cost for inserting or deleting a node or edge is set to $1.0$.
Node Substitution
The substitution cost $c_{sub}(n_{1},n_{2})$ between two nodes is determined by:
$c_{sub}(n_{1},n_{2})=\begin{cases}0.0&\text{if }\text{label}(n_{1})=\text{label}(n_{2})\land\text{type}(n_{1})=\text{type}(n_{2})\\ 0.5&\text{if }\text{label}(n_{1})=\text{label}(n_{2})\land\text{type}(n_{1})\neq\text{type}(n_{2})\\ 1.0&\text{otherwise}\end{cases}$ (1)
Edge Matching
Edges are matched based on the equality of their normalized labels.
The partial penalty of $0.5$ for type mismatches allows us to distinguish cases where the model correctly inferred the intent (the label) but selected the wrong BPMN element type (e.g., using a generic Task instead of a User Task) from cases where the model hallucinated an entirely incorrect step.
RGED normalizes the GED by considering the complexity of the involved graphs, providing a metric that is independent of diagram size:
$\text{Relative GED}=\frac{\text{GED}(G_{1},G_{2})}{\text{GED}(G_{1},\emptyset)+\text{GED}(G_{2},\emptyset)}$ (2)
@@
where $\emptyset$ represents an empty graph. This results in a similarity score where $1.0$ represents a semantically identical structure and $0.0$ represents total dissimilarity.
To implement the calculation, we utilized the NetworkX Python library, which provides efficient algorithms for computing graph edit distance. Unlike prior approaches that rely on domain-specific heuristics or ML-based enhancements [39], our implementation applies standard GED operations to measure similarity between BPMN diagrams, albeit with the pre-processing step of semantic normalization.
By minimizing the edit distance to human-generated ground truth models, we utilize GED and RGED as proxies for structural fidelity and logical correctness. However, it is important to note that these metrics do not explicitly capture higher-level qualitative factors such as visual readability, layout simplicity, or cognitive load, which are emphasized in studies like those by Huang and Kumar [17] and Pavlicek et al. [33].
@@ 4.3 Comparison with Baseline XML Editing
To establish the effectiveness of BPMN Assistant, we conducted a comparative analysis between our atomic editing approach and a full-regeneration baseline (implemented here as direct XML generation). This baseline serves as a proxy for conversational refinement workflows used by state-of-the-art tools such as BPMN-Chatbot [20] and ProMoAI [23]. Both systems prompt the LLM to produce a revised structured representation of the process conditioned on the current model state (e.g., intermediate JSON in BPMN-Chatbot, or constrained code/POWL construction in ProMoAI). BPMN-Chatbot explicitly targets token efficiency via its intermediate JSON representation, whereas ProMoAI primarily leverages formal structure and constrained generation to improve reliability.
For the XML baseline, we utilized a system prompt instructing the model to act as a BPMN expert and output valid BPMN 2.0 XML directly, without intermediate steps or function calling. This represents the standard ’zero-shot’ approach common in current LLM applications. For this comparison, we selected a representative set of diagram modification tasks and executed them using both representations. We evaluated the approaches using RGED and its derived similarity score, as defined in Equations (2) and (3).
@@ 4.4 Generation Accuracy
For the purposes of this evaluation, a "failure" is defined as any generation event that resulted in an unrenderable model. This encompasses both syntactic invalidity—such as unclosed XML tags or malformed JSON objects that prevent parsing—and structural inconsistencies, specifically reference hallucinations where sequence flows target node identifiers that do not exist within the element set. Any output falling into these categories was recorded as a failure in Tables 4 and 5 and treated as a non-functional result.
The comparative analysis revealed that our JSON-based representation achieved an average similarity score of 0.72 compared to 0.70 for direct XML generation. While this indicates a slight numerical advantage for JSON, the difference is minimal, suggesting that both representations perform similarly in terms of structural accuracy. However, JSON demonstrated greater reliability, with fewer total failures across models.
@@
Beyond similarity scores, JSON continues to outperform XML in terms of efficiency, a trend that holds even with newer, faster models. Latency measurements represent the total API response time recorded via the respective commercial providers (OpenAI, Anthropic, Google, and Fireworks AI). This metric reflects the real-world performance experienced by end-users of cloud-based LLM services. As shown in Table 6, JSON-based BPMN generation achieved a mean latency of 13.42 seconds compared to 24.82 seconds for XML. Additionally, while JSON requires a richer input prompt (averaging 2,678 tokens), it produces significantly more concise outputs (688 tokens vs 1,832 tokens). This reduction in output tokens is particularly advantageous for cost scaling, as output tokens are typically significantly more expensive than input tokens. Furthermore, the financial overhead of the increased input context is effectively mitigated by prompt caching technologies, which drastically reduce the cost of processing the static schema definitions and instructions required by our approach.
@@
Mean Latency (seconds) 13.42 24.82
Average Input Tokens 2678.63 474.05
Average Output Tokens 688.04 1832.48
@@ 4.5 Editing Capabilities
In addition to diagram generation, we evaluated the models’ ability to interpret natural language editing instructions and perform the requested modifications. For this evaluation phase, we manually curated a dataset of 40 specific modification requests targeting various process elements. We employed a binary success/fail metric, as editing tasks involve correctly understanding the request and applying appropriate modifications to an existing diagram without corrupting the remaining structure.
To determine validity, we employed a two-stage verification process. First, a Python-based automated validator checked all outputs for syntactic correctness and referential integrity (e.g., ensuring no broken XML tags or references to non-existent IDs). Outputs failing this stage were automatically classified as failures. Second, the syntactically valid outputs underwent expert manual verification to assess semantic correctness. This step involved a binary check against the editing prompt (e.g., verifying that a "delete task" operation actually removed the target node and correctly reconnected the surrounding sequence flows). Given the objective nature of these boolean operations, a single expert evaluator was deemed sufficient.
The results, summarized in Table 7, show that models consistently achieved higher success rates when using our JSON-based approach compared to direct XML manipulation. The performance gap is particularly evident in open-weight models; for instance, DeepSeek V3 achieved a 50% success rate with JSON but failed almost entirely with XML (8%). This suggests that the structured intermediate representation helps mitigate the complexities of verbose XML syntax, allowing models to focus on the logical changes rather than the syntactic overhead of the BPMN standard.
Even for flagship models like Claude 4.5 Sonnet, which achieved an 85% success rate in both modalities, the JSON approach offers superior programmatic verifiability. Because the JSON output adheres to a strict internal schema, the system can validate logical consistency (e.g., ensuring all branch targets exist) before conversion. In contrast, validating generated XML often requires complex parsing that may fail on minor syntax hallucinations.
@@
As detailed in Table 8, the trade-off for this reliability is context size. The JSON editing approach requires supplying the full BPMN intermediate representation and various process examples in the prompt, resulting in an average of 22,071 input tokens compared to 5,149 for XML. However, this ’up-front’ cost pays dividends in speed: the JSON approach reduces generation latency by nearly 43% (20.35s vs 35.63s) and reduces output verbosity by over 75% (607 vs 2,630 tokens). In an interactive tool, this lower latency is critical for user experience.
@@
Average Latency (s) 20.35 35.63
Average Input Tokens 22,071.44 5,149.42
Average Output Tokens 607.92 2,630.44
@@ 5 Discussion
These findings empirically validate the initial conceptual assumption of this work, namely that constraining LLM interaction through structured representations leads to more reliable and controllable BPMN model manipulation.
5.1 The Efficacy of Structured Representations (RQ1)
Our evaluation reveals a distinct dichotomy between generation and editing performance. In de novo generation tasks, the difference between JSON and XML approaches was marginal (e.g., GPT-5.1 achieved 0.73 with JSON vs 0.70 with XML). This suggests that for "tabula rasa" tasks, modern LLMs are sufficiently capable of managing XML syntax when generating from scratch.
However, the editing tasks expose the fragility of direct XML manipulation. The JSON-based approach achieved consistently higher success rates across all models. We attribute this difference to the verbosity and structural fragility of direct XML manipulation. When editing XML, an LLM must navigate verbose opening and closing tags, often losing track of the hierarchical structure or hallucinating invalid ID references. In contrast, our JSON schema and the associated atomic editing functions (e.g., add_element) force the model to focus on the logical operation rather than the syntactic overhead. This structural constraint effectively acts as a guardrail, ensuring that modifications remain valid by design.
5.2 Democratizing Process Modeling (RQ2)
Perhaps the most significant finding is the impact of our approach on open-weights models. While frontier proprietary models like Claude 4.5 Sonnet performed admirably in XML editing (85%), open-weights models such as DeepSeek V3 struggled significantly, achieving only an 8% success rate with XML. However, when switching to our JSON-based function calling approach, DeepSeek V3’s success rate jumped to 50%.
This has profound implications for enterprise adoption. Many organizations in regulated industries (finance, healthcare) cannot send sensitive process data to external APIs like OpenAI or Anthropic due to data privacy concerns. Our results demonstrate that by using a structured intermediate representation, organizations can deploy locally hosted, open-weights models to perform complex process modeling tasks that were previously the domain of massive frontier models. This effectively lowers the barrier to entry for secure, on-premise AI process assistants.
5.3 Efficiency and Latency Trade-offs (RQ3)
The shift to a JSON-based editing workflow introduces a counter-intuitive trade-off: we drastically increase the input context size to decrease latency. As shown in Table 8, the JSON approach requires supplying the entire current state of the process, leading to a $\sim 4\times$ increase in input tokens compared to the XML baseline.
However, in the current landscape of LLM economics and performance, this trade-off is highly favorable. Input tokens are computationally cheaper and faster to process than output tokens. By shifting the complexity to the input (the context) and restricting the output to concise JSON function calls, we achieved a 43% reduction in total latency (20.35s vs 35.63s). For interactive tools where user experience depends on responsiveness, this reduction is critical. Furthermore, the 75% reduction in output tokens leads to significant cost savings at scale, as output tokens typically cost 3-4 times more than input tokens across major providers. This efficiency is further amplified by prompt caching, which allows the static JSON schema definitions—the bulk of our input context—to be processed at a fraction of the standard input cost.
5.4 Comparison with Existing Approaches
Unlike previous tools such as BPMN-Chatbot [20] and ProMoAI [23], which rely on regenerating the full process definition (via JSON or code) for every refinement, our system separates the editing logic from the model generation. While ProMoAI leverages formal methods (POWL) and BPMN-Chatbot utilizes intermediate JSON to optimize model creation, both approaches fundamentally treat modification as a conversational regeneration task. In contrast, BPMN Assistant implements a dedicated intermediate layer that supports atomic function calling, enabling targeted and deterministic updates without the overhead or instability associated with regenerating the entire process state for minor edits.
@@ 6 Limitations and Future Work
BPMN Assistant, while demonstrating promising results, has limitations that should be acknowledged. Although the system supports a robust set of elements, including intermediate events (timer, message) and inclusive gateways, the current implementation does not support collaboration diagrams (pools and lanes) or complex artifacts like data objects. This restricts the system’s application in scenarios requiring multi-participant modeling. However, as the focus of this work is on executable process orchestration, this omission does not impact the primary research objectives.
@@
The performance of BPMN Assistant is heavily dependent on the underlying LLM. Our evaluation revealed significant variations in performance across different models, with Claude 4.5 Sonnet and GPT-5.1 showing particularly strong results. Further investigation into the specific capabilities and limitations of different LLMs in the context of process modeling would provide valuable insights for future improvements.
@@ 7 Conclusion
This paper presented BPMN Assistant, a novel approach to business process modeling that leverages LLMs to bridge the gap between natural language descriptions and formal BPMN representations. Through a comprehensive evaluation using state-of-the-art models, including GPT-5.1, Claude 4.5 Sonnet, and DeepSeek V3, we demonstrated that using a simplified JSON structure instead of raw XML yields significant benefits in reliability and efficiency.
Our results indicate that while direct XML generation is feasible for creating simple models from scratch, it is often insufficient for iterative modification tasks. The proposed JSON-based approach outperformed or matched direct XML manipulation across all tested models in editing scenarios. While frontier models like Claude 4.5 Sonnet achieved parity in both modalities, the JSON advantage was critical for other models. This was most pronounced in open-weights models, where the structured approach enabled DeepSeek V3 to achieve a viable 50% success rate compared to a near-total failure (8%) with standard XML prompting. This suggests that intermediate representations are a key enabler for deploying cost-effective, locally hosted models in process automation workflows.
In terms of performance, the trade-off between increased input context and reduced output complexity proved highly advantageous. By shifting the computational burden to the input context, our approach reduced editing latency by approximately 43% and decreased output token usage by over 75%. These efficiency gains are critical for the practical adoption of LLM-based tools in interactive, real-time environments.
Looking ahead, future work will focus on conducting comprehensive usability studies to validate the system’s effectiveness with non-technical business users. Additionally, we aim to further optimize the JSON schema to better handle complex cyclic dependencies and improve the system’s robustness when handling highly ambiguous natural language descriptions.
@@ Appendix A Appendix: Process Example
The following examples demonstrate how the JSON intermediate representation handles complex BPMN constructs, including cyclic flows (loops) and event definitions.
Example 1: Cyclic Flow (Loops)
This example demonstrates the use of the next attribute to model a loop where a student must retake an exam upon failure. This illustrates the system’s ability to handle non-hierarchical flows.
@@
"Someone starts a process by entering an exam room. After that, they take the test. Once the exam is finished, their score is checked. If they scored more than 50%, their grade is recorded and the process ends. But if they fail, they have to go back to the beginning and take the exam again."
JSON Representation:
    "process": [
        "type": "startEvent",
        "id": "start"
        "type": "task",
        "id": "task1",
        "label": "Enter exam room"
        "type": "task",
        "id": "task2",
        "label": "Take test"
        "type": "exclusiveGateway",
        "id": "eg",
        "label": "Score more than 50%?",
        "has_join": false,
        "branches": [
            "condition": "Yes",
            "path": [
                "type": "task",
                "id": "task3",
                "label": "Enter grade"
                "type": "endEvent",
                "id": "end1"
            "condition": "No",
            "path": [],
            "next": "task1" // Redirects flow back to the first task
Example 2: Event-Driven Workflow
This example demonstrates the support for specific event definitions (Timer and Message events) required for executable process models.
Textual description:
"A batch process starts at midnight using a timer. The system processes pending orders and waits for 5 minutes. Then it sends a completion notification and waits to receive a confirmation message before ending."
@@
      "label": "Start at midnight",
      "eventDefinition": "timerEventDefinition"
@@
      "label": "Process pending orders"
      "type": "intermediateCatchEvent",
      "id": "timer1",
      "label": "Wait 5 minutes",
      "eventDefinition": "timerEventDefinition"
      "type": "intermediateThrowEvent",
      "id": "message1",
      "label": "Send completion notification",
      "eventDefinition": "messageEventDefinition"
      "type": "intermediateCatchEvent",
      "id": "message2",
      "label": "Wait for confirmation",
      "eventDefinition": "messageEventDefinition"